OpenStudy (anonymous):

What is the solution to the 2 equations? 3x - 5y = 11 2x + 4y = 9

OpenStudy (anonymous):

so technically r both equations x and y the same number?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is hard. just plug in numbers that equal to it

OpenStudy (whpalmer4):

Yes, the solution to the system of equations will have the same value for \(x\) in both, and the same value for \(y\) in both.

OpenStudy (anonymous):

i will leave then

OpenStudy (anonymous):

I have to graph them and then find the solution but Im having a hard time graphing it

OpenStudy (whpalmer4):

If you graph both equations, you will get two straight lines (though one may be on top of the other), and if they intersect, the point of intersection is the solution.

OpenStudy (mathmale):

@yuna123, I know you'd like to help, and appreciate that, but your comments so far have been off the wall. Please suggest methods such as substitution, elimination, graphing, and so on, methods that truly succeed.

OpenStudy (whpalmer4):

My easy approach to graphing lines: put \(x=0\) solve for value of \(y\). Put a point at \(0,y\). Put \(y=0\), solve for value of \(x\). Point a point at \(x,0\). Draw a straight line through the points. Done.

OpenStudy (anonymous):

My teacher has only given me a graph that shows quadrant 1

OpenStudy (whpalmer4):

Does it show the two lines?

OpenStudy (whpalmer4):

Ah, no, I guess you are supposed to use that as graph paper.

OpenStudy (anonymous):

My question says Write the equations in slope intercept form. Graph the pair of linear equations. Use the graph to estimate the solution to the system of equations.

OpenStudy (whpalmer4):

Okay, helpful if you provide all the info up front. Slope-intercept form is \[y = mx + b\]where \(m\) is the slope and \(b\) is the \(y\)-intercept. To put your equations in this form, just solve them for \(y\) alone on the left side of the \(=\).

OpenStudy (anonymous):

3x - 5y = 11 would be y = 0.6x - 2.2? and 2x + 4y = 9 would be y = -0.5x + 2.25?

OpenStudy (whpalmer4):

I would keep them as fractions, personally, but those appear correct.

OpenStudy (anonymous):

So then they would be y = -2/4x + 2.25? y = -3/5x - 2.2?

OpenStudy (anonymous):

There is no solution then...

OpenStudy (denisaboichuk):

y=25

OpenStudy (whpalmer4):

No, I wouldn't use any decimals whatsoever. \[3x - 5y = 11\]\[ 2x + 4y = 9\] \[3x-5y = 11\]\[5y = 3x-11\]\[y = \frac{3}{5}x-\frac{11}{5}\] \[2x+4y=9\]\[4y = -2x+9\]\[y = -\frac{1}{2}x+\frac{9}4\] Those lines have different slopes, so there most definitely IS a solution. @denisaboichuk that is not it.